OpenStudy (bossimbacon):

Complete the standard form of the equation . 3y/2=8-x ___ x +___y = 16

OpenStudy (bossimbacon):

@StudyGurl14

OpenStudy (anonymous):

3y/2=8-x 3y=2(8-x) 3y=16-2x 2x+3y=16